OpenStudy (anonymous):

Funny math question..Dare to try? (not a trick question)? 7 girls on a bus being driven to school each girl has 7 bags for each bag there are 7 big cats for every big cat there are 7 little cats how many legs are on the bus? I know the answer:) do u?

OpenStudy (anonymous):

7*2=14 7*4=28 7*4=28 add = 70

OpenStudy (anonymous):

nope :)

OpenStudy (anonymous):

no

OpenStudy (anonymous):

14

OpenStudy (anonymous):

14 right ans

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

16

OpenStudy (anonymous):

1568

OpenStudy (hba):

The bus has wheels.Therfore 0 legs :D kthanksbye.

OpenStudy (anonymous):

nope :)

OpenStudy (anonymous):

16 + 343x4 + 2401x4

OpenStudy (anonymous):

406

OpenStudy (hba):

As a math question, 10990 legs altogether.

OpenStudy (anonymous):

Him1618 is right :D

OpenStudy (anonymous):

10992 4 legs x 7 little x 7 big x 7 bags x 7 girls 4 legs x 7 big x 7 bags x 7 girls 2 legs x 7 girls 2 legs for the driver.

OpenStudy (hba):

@Falkqwer We dont have to count the legs of the driver :)

OpenStudy (hba):

Please see your question :)

OpenStudy (anonymous):

YES u do lol

OpenStudy (anonymous):

but question is on the bus

OpenStudy (hba):

what about the conductor then 10994 legs

OpenStudy (anonymous):

right ans is 16

OpenStudy (anonymous):

7 girls on a bus being DRIVEN to school

OpenStudy (anonymous):

*shoots self*

OpenStudy (hba):

I bet there is a conductor also so the answer is 10994 legs

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

lol soooo much hate... :'( why cant u guys just be nice?

OpenStudy (hba):

*Seriously shoot yourself*

OpenStudy (anonymous):

Fine *unshoots self*

OpenStudy (anonymous):

i think its funny :D

OpenStudy (hba):

Lol ^

OpenStudy (anonymous):

:D Well thx to everyone who tried :D

OpenStudy (hba):

Whatever.

OpenStudy (anonymous):

i would have but i saw the awnser

OpenStudy (anonymous):

Thanks for the laugh...I was bored!!

OpenStudy (anonymous):

nvm lol too late :P

OpenStudy (anonymous):

XD np Good try tho :)

OpenStudy (hba):

Thanks for playing Riddles.

OpenStudy (anonymous):

Well...thx everyone for coming ^_^

OpenStudy (hba):

You are quiet welcome.